Rich is twice as old as chris. Five years from now the sum of their ages will be 52. How old is chris?
Answer by lwsshak3(11628)

Rich is twice as old as chris. Five years from now the sum of their ages will be 52. How old is chris?
let x=chris' age
2x=rich's age
..
5yrs from now:
x+5=chris' age
2x+5=rich's age
..
(x+5)+(2x+5)=52
3x+10=52
3x=42
x=14
2x=28
chris' age=14